Perpetual Inventory Using FIFO

Beginning inventory, purchases, and sales for Item B901 are as follows:

May 1 Inventory 39 units @ $19
9 Sale 30 units
13 Purchase 35 units @ $20
28 Sale 17 units

Assuming a perpetual inventory system and using the first-in, first-out (FIFO) method, determine (a) the cost of merchandise sold on May 28 and (b) the inventory on May 31.

a. Cost of merchandise sold on May 28 $
b. Inventory on May 31
Answer #1

Under FIFO ,units acquired first are sold first.Also under perpetual inventory system ,inventory records are maintained date wise.

FIFO METHOD
PURCHASE COST OF GOODS SOLD INVENTORY
Quantity unit cost Total Quantity unit cost Total Quantity unit cost Total
May 1 39 19 741
9 30 19 570 39-30=9 19 171
13 35 20 700 9 19 171
35 20 700
28 9 19 171
8 20 160 35-8= 27 20 540

May 28 sale of 17 units are made as 9 unit at the rate of $ 19 and remaining 8 units are sold at the rate of $ 20.

A)cost of merchandise sold on May 28 : 171 +160 = 331

B) Inventory on May 3 1 : 27 units at the unit cost of $ 20 = $ 540